OpenStudy (anonymous):

Find -f(x) pf f(x) = [/sqrt(x+3)]/

OpenStudy (anonymous):

\[f(x) = \sqrt{x-3}\] ***

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (cruffo):

You want me to give it a shot?

OpenStudy (ash2326):

You need to find f(-x) or - f(x) ??

OpenStudy (anonymous):

-f(x) and yes please give it a shot/

OpenStudy (cruffo):

\[-f(x) = -1\cdot f(x)\]

OpenStudy (cruffo):

\[ = -\sqrt{x-3}\] done

OpenStudy (ash2326):

Yes just multiply both sides by -1

OpenStudy (anonymous):

oh sick, thanks guys :)